OpenStudy (anonymous):

anyone know trigonometry by solving equations such as cot(3x)(cot(x) - 1) = 0

OpenStudy (faiqraees):

okay do you know trigonometric identities?

OpenStudy (anonymous):

either cot 3x=0 \[\cot3x=0=\cot \left( 2n+1 \right)\frac{ \pi }{ 2 },n~is~an~integer.\] \[3x=\left( 2n+1 \right) \frac{ \pi }{ 2 },x=?\] or cotx-1=0 \[\cot x=1=\cot \left( n \pi+\frac{ \pi }{ 4 } \right),n~is~an~integer.\] x=?

OpenStudy (faiqraees):

or more simplified way \[\cot{3x}=0\]\[1/tan{3x}=0\]\[1=tan{3x}*0\]\[no~possible~solutions\] \[\cot{x}-1=0\]\[\cot{x}=1\]\[\ 1/tan{x}=1\]\[\tan{x}=1\]\[\ x=tan^{-1}1\]\[x=45,225\]